OpenStudy (anonymous):

Hello, could someone help me with this? a boat traveled 140miles downstream and back. The trip downstream took 7hours. the trip back took 70hours. What is the speed of the boat in still water? What is the speed of the current?

OpenStudy (anonymous):

w= speed of water v= speed of boat

OpenStudy (anonymous):

(w+v)7=140 (w-v)70=140

OpenStudy (anonymous):

Solve for w and v

OpenStudy (anonymous):

how do i solve the w and y??

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

(w+v)7=140 (w-v)7=14 7w+7v=140 7w-7v=14 ------------- 14w=154 w=11 (11-v)7=14 11-v=2 v=9

OpenStudy (anonymous):

thank you soo much!

OpenStudy (anonymous):

So i wasnt that wrong with my first answer haha!

OpenStudy (anonymous):

where is your first answer?

OpenStudy (anonymous):

ive tried the first part but wasnt very sure on how to follow the steps as you did.